OpenStudy (anonymous):

If the cos 25° is 0.906, then the sin of x° is 0.906. Find x.

OpenStudy (anonymous):

@sayakdbz

OpenStudy (anonymous):

a'ight. hold on

OpenStudy (anonymous):

Okay and tell me how you got the answer cause I have nooo idea how to do this one

OpenStudy (anonymous):

that should be approximately 65 degrees :3

OpenStudy (***[isuru]***):

\[\sin ( 90 - \theta ) = \cos \theta\]

OpenStudy (***[isuru]***):

it's the theory used here

OpenStudy (anonymous):

just what @***[ISURU]*** said. guess i wasn't fast enough. give him a medal ^_^

OpenStudy (***[isuru]***):

\[\cos \ 25 = \sin ( 90 - 25 ) = \sin ( 65)\]

OpenStudy (anonymous):

The distance is the adjacent side of the 65 degree angle and the opposite is 25 degrees.